OpenStudy (anonymous):

I have a first order differential equation (dq/dt) + 100q = 0.2, How do I find the integrating factor here?

OpenStudy (asnaseer):

this is already in standard form:\[q'+p(t)q=q(t)\]where p(t)=100 and q(t)=0.2 the integrating factor is found by:\[u(t)=e^{\int p(t)dt}\]

OpenStudy (anonymous):

if the equation is of form y'+yP(x)= Q(x) then Integrating factor = \[e ^{\int\limits_{?}^{?}P(x)dx}\] so here in your equation integrating factor = e^(100x)